# Revisiting the operator extension of strong subadditivity

### v2

@@ Abstract
It was observed in [1, Sec. 3] that (1) resembles—but does not follow from—the monotonicity of the relative modular operator, which takes the form $\rho\otimes\sigma^{-1}$ for two density matrices $\rho,\sigma$. In the following, we will see that (1) is an immediate consequence of Connes’ theory of spatial derivatives on von Neumann algebras [4], which can be understood as a generalization of (relative) modular theory to settings where a given von Neumann algebra $M\subset B({\mathcal{H}})$ is not in standard form, i.e., there is no cyclic and separating vector $\Omega\in{\mathcal{H}}$ for $M$. Indeed, (1) is simply the monotonicity of the spatial derivative for inclusions of matrix factors. This observation has two consequences:
@@
    There exists an appropriate generalization of (1) to the setting of von Neumann algebras. In the standard representation, it reduces to the monotonicity of the relative modular operator [5, Eq. (II.1.3)].
@@ 2 General case
A vector $\xi$ is called $\omega$-bounded if $R^{\omega}(\xi)$ is a bounded operator, i.e., there is a constant $C_{\xi}>0$ such that $\lVert a\xi\rVert\leq C_{\xi}\,\omega(a^{*}a)^{1/2}$ for all $a\in A$. The set of $\omega$-bounded vectors is a dense subspace in $\mathcal{H}$} [7, Lemma 11.2]. If $\xi$ is $\omega$-bounded we denote the bounded extension to an operator $L^{2}(A,\omega)\to{\mathcal{H}}$ again by $R^{\omega}(\xi)$, and we set
@@
This concludes the proof of Lemma˜1.
@@
We have deduced (11) from (14). As a final remark, we note that the two inequalities are, in fact, equivalent. Indeed, in the setting of Lemma˜1, Eq.˜11 implies

### v3

@@ Abstract
It was observed in [1, Sec. 3] that (1) resembles—but does not follow from—the monotonicity of the relative modular operator, which takes the form $\rho\otimes\sigma^{-1}$ for two density matrices $\rho,\sigma$. In the following, we will see that (1) is an immediate consequence of Connes’ theory of spatial derivatives on von Neumann algebras [4], which can be understood as a generalization of (relative) modular theory to settings where a given von Neumann algebra $M\subset B({\mathcal{H}})$ is not in standard form, i.e., there is no cyclic and separating vector $\Omega\in{\mathcal{H}}$ for $M$. Indeed, (1) is simply the well-known monotonicity of the spatial derivative [5, Lem. 4.2], applied to the inclusions of matrix factors. This observation has two consequences:
@@
    There exists an appropriate generalization of (1) to the setting of von Neumann algebras. In the standard representation, it reduces to the monotonicity of the relative modular operator [6, Eq. (II.1.3)].
@@ 2 General case
A vector $\xi$ is called $\omega$-bounded if $R^{\omega}(\xi)$ is a bounded operator, i.e., there is a constant $C_{\xi}>0$ such that $\lVert a\xi\rVert\leq C_{\xi}\,\omega(a^{*}a)^{1/2}$ for all $a\in A$. The set of $\omega$-bounded vectors is a dense subspace in $\mathcal{H}$} [8, Lemma 11.2]. If $\xi$ is $\omega$-bounded we denote the bounded extension to an operator $L^{2}(A,\omega)\to{\mathcal{H}}$ again by $R^{\omega}(\xi)$, and we set
@@
This concludes the proof of Lemma˜1. We note that Eq.˜14 remains valid without assuming the faithfulness of $\omega$ (see [5, Lem. 4.2] for further details).
@@
We have deduced (11) from (14). As a final remark, we note that the two inequalities are, in fact, equivalent. Indeed, in the setting of Lemma˜1, Eq.˜11 implies (cf. [5, Eq. (9.1)])